Kaitlin invested her savings in two investment funds. The amount she invested in Fund A was twice as much as the amount she invested in Fund B. Fund A
returned a 5% profit and Fund B returned a 3% profit. How much did she invest in Fund B, if the total profit from the two funds together was $1040?

Answer by ikleyn(52781)

.
Kaitlin invested her savings in two investment funds.
The amount she invested in Fund A was twice as much as the amount she invested in Fund B.
Fund A returned a 5% profit and Fund B returned a 3% profit.
How much did she invest in Fund B, if the total profit from the two funds together was $1040?
~~~~~~~~~~~~~~~~~~

The amount invested in Fund B = x dollars;

The amount invested in Fund A = 2x dollars.


Write the total interest equation

    0.05*(2x) + 0.03*x = 1040  dollars.


Simplify and find x

    0.1x + 0.03x = 1040

       0.13x     = 1040

           x     = 1040/0.13 = 8000.


ANSWER.  $8000 invested in Fund B.

Solved.

-------------

In this problem, a professional term is "the interest", or "the annual interest", instead of "profit".

The term "profit" has different meaning and stands for another concept.

---------------

It is a standard and typical problem on investments.

If you need more details,  or if you want to see other similar problems solved by different methods,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.

You will find there different approaches  (using one equation or a system of two equations in two unknowns),  as well as
different methods of solution to the equations  (Substitution,  Elimination).

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.